OpenStudy (anonymous):

During a race, a runner runs at a speed of 6 m/s. 2 seconds later, she is running at a speed of 10 m/s. What is the runner’s acceleration? Show your work

OpenStudy (anonymous):

change velocity divided by change in time

OpenStudy (anonymous):

(10-6)m/s/2s=2m/s^2

OpenStudy (anonymous):

Hope that's clear

OpenStudy (anonymous):

I HOPE SO TO..LOL

OpenStudy (anonymous):

a = (v-u)/t =(10-6)/2 =4/2 =2m/s^2